OpenStudy (anonymous):

Which of the following is not a possible value for a probability? A. 95% B. 0.05 C. .999999999 D. 0 E. 11.05

OpenStudy (anonymous):

E. Probability can't be greater than 1 or 100%

OpenStudy (anonymous):

idk , this help seriously !!

OpenStudy (amistre64):

tag!! running a test

OpenStudy (anonymous):

uqqh .

OpenStudy (anonymous):

The sum of all the probabilities in a sample space must equal 1, and none of them can have negative possibility (that doesn't make sense). Therefore none of them can have a probability greater than 1.

OpenStudy (anonymous):

If two dice are rolled, what is the probability of getting two numbers greater than 4?

OpenStudy (amistre64):

4out of 12

OpenStudy (anonymous):

wouldnt it be 1/36

OpenStudy (amistre64):

2 numbers added higher than 4? or 2 faces that are higher than 4?

OpenStudy (anonymous):

What is the probability for 1 die getting a number greater than 4?

OpenStudy (amistre64):

2 numbers implies die faces I think

OpenStudy (anonymous):

Assume fair 6 sided dice. Then the numbers greater than 4 are 5 and 6. Since dice are independent. P(A and B)=P(A)*P(B)

OpenStudy (anonymous):

2/6 = 1/3. To get 2 dice with each having a number greater than 4 you'd multiply the probability of each one together by the product rule.

OpenStudy (anonymous):

1/3

OpenStudy (amistre64):

2 faces out of 6 are higher than 4; so if I see it right

OpenStudy (anonymous):

Two independent outcome occurring here so as Xavier said you have P(A and B) = P(A)*P(B)

OpenStudy (anonymous):

Right? If each one has a 1/3 chance, having two at the same time would be \(\frac{1}{3}*\frac{1}{3} = \frac{1}{9}\)

OpenStudy (amistre64):

tag

OpenStudy (anonymous):

Seriously.. it's just annoying.